The cable TV industry reports that 59% of homes in the US to cable TV. If you pick four US households at random, what is the probability that at least one of the households is not a cable TV subscriber?
Answer by stanbon(75887)

The cable TV industry reports that 59% of homes in the US to cable TV. If you pick four US households at random, what is the probability that at least one of the households is not a cable TV subscriber?
------------------------
Binomial Problem:
n = 4 ; p = 0.59 ; q = 0.41 ; 1<= x <=4
---------------------------------------------
P(at least one) = 1 - P(none)
---
P(at least DISABLED_event_one= 1 - 4C0(0.59)^0*(0.41)^4 = 1-0.41^4 = 0.9717...
Cheers,
Stan H.